# SubTrack your Grad: Gradient Subspace Tracking for Memory and Time Efficient Full-Parameter LLM Training

### v1

@@ 6 Discussion and Conclusion
We proposed a time- and memory-efficient method that projects gradients into a lower-dimensional subspace, and preserves the previously computed subspace to incorporate gradient components from the orthogonal complement to perform rank-$1$ updates. This approach reduces the frequency of abrupt transitions between iterations and leverages the available information effectively.
In some cases, the extent of changes in the subspace may require updates of ranks greater than $1$; the pre-training of Llama-based architectures is a clear example of this need. During our experiments, we observed that applying updates as per (6), using the full-rank SVD of the tangent vector from (5), can hinder convergence if the singular values of the tangent vector become very small. Furthermore, simultaneously updating vector spaces associated with different singular values caused convergence issues in some cases. Therefore, we restricted updates to rank-$1$ in this paper, as this approach still enabled SubTrack-Grad to achieve performance comparable to or better than GaLore. In future work, we plan to explore increasing the rank of updates without compromising convergence. Additionally, dynamically selecting the step size could eliminate the need for manual tuning as a hyperparameter which is also left for future works.
Surprisingly, SubTrack-Grad outperforms full-rank fine-tuning on certain tasks, suggesting that low-rank optimization may also serve as a regularizer. Exploring this property further is also an exciting direction for future work.
Appendix A Convergence of SubTrack-Grad
@@
proof. To demonstrate that SubTrack-Grad converges to the global minimum during training, we begin by deriving the recursive form of the gradients.
@@
While SubTrack-Grad utilizes right/left projections to reduce memory consumption, the proof is presented using both projection matrices to ensure generality. Here, we demonstrate how the proof proceeds under the assumption $m\leq n$ (without loss of generality), which allows the use of the left projection matrix.
@@ Appendix B Grassmann Exponential
$\Delta=Q\begin{pmatrix}0&B^{T}\\ B&0\end{pmatrix}Q^{T}=\left[Q\begin{pmatrix}0&-B^{T}\\ B&0\end{pmatrix}Q^{T},P\right]$ (28)
@@
$\Delta^{\text{hor}}_{Q}=[\Delta,P]Q=Q\begin{pmatrix}0&-B^{T}\\ B&0\end{pmatrix}$ (29)
@@
The orthogonal group $O(n)$ is defined as the set of all $n\times n$ matrices $Q$ over $\mathbb{R}$ such that $Q^{T}Q=QQ^{T}=I_{n}$, where $Q^{T}$ is the transpose of $Q$ and $I_{n}$ is the $n\times n$ identity matrix:
$O(n)=\{Q\in\mathbb{R}^{n\times n}\mid Q^{T}Q=I_{n}=QQ^{T}\}.$
Then the tangent space of the orthogonal group $O(n)$ at a point $Q$, denoted $T_{Q}O(n)$, is defined as the set of matrices of the form $Q\Omega$, where $\Omega\in\mathbb{R}^{n\times n}$ is a skew-symmetric matrix, i.e., $\Omega^{T}=-\Omega$:
$T_{Q}O(n)=\{Q\Omega\mid\Omega\in\mathbb{R}^{n\times n},\Omega^{T}=-\Omega\}.$
@@
If $P\in Gr(n,p)$ and $\Delta\in T_{P}Gr(n,p)$ with $\Delta^{\text{hor}}_{Q}=Q\begin{pmatrix}0&-B^{T}\\ B&0\end{pmatrix}$, the geodesic in the Grassmannian is therefore
$\text{Exp}^{Gr}_{P}(t\Delta)=\pi^{OG}\left(Q\exp_{m}\left(t\begin{pmatrix}0&-B^{T}\\ B&0\end{pmatrix}\right)\right).$ (31)
@@
$B=U_{\perp}^{T}\Delta^{\text{hor}}_{U}=U_{\perp}^{T}\hat{Q}\Sigma V^{T}$
@@
with $W:=U_{\perp}^{T}\hat{Q}\in St(n-p,r),\Sigma\in\mathbb{R}^{r\times r},V\in St(p,r)$. Let $W_{\perp},V_{\perp}$ be suitable orthogonal completions. Then,
$\exp_{m}\begin{pmatrix}0&-B^{T}\\ B&0\end{pmatrix}=\begin{pmatrix}V&V_{\perp}&0&0\\ 0&0&W&W_{\perp}\end{pmatrix}\begin{pmatrix}\cos(\Sigma)&0&-\sin(\Sigma)&0\\ 0&I_{p-r}&0&0\\ \sin(\Sigma)&0&\cos(\Sigma)&0\\ 0&0&0&I_{n-p-r}\end{pmatrix}\begin{pmatrix}V^{T}&0\\ V_{\perp}^{T}&0\\ 0&W^{T}\\ 0&W_{\perp}^{T}\end{pmatrix},$
@@
which leads to the desired result when inserted into (31). For more mathematical details, you can refer to Edelman et al. (1998), Bendokat et al. (2024), or other useful resources on Grassmann geometry.
Appendix C Time Complexity Analysis
@@
Table 5 presents the time complexity breakdown for the subspace update step in the SubTrack-Grad algorithm assuming a $m\times n$ gradient matrix and rank $r$ projection matrix, where $r\ll m\leq n$. As outlined in Algorithm 1, the subspace update step begins by solving the least squares problem (3) to estimate the optimal update for $S_{t}$, the $m\times r$ orthonormal matrix. This operation has a time complexity of $O(mr^{2})$. Computing the residual and the partial derivative with respect to $S_{t}$ requires $O(mrn)$ and $O(mnr)$ time respectively. This is because the solution to the least squares problem, $A$, has shape $r\times n$ which is multiplied by $S_{t}$ in the residual $R=G_{t}-S_{t}A$, resulting in time complexity $O(mrn)$. The following operation for the partial derivative is $-2RA^{T}$, where the matrix multiplication has $O(mnr)$ complexity. The tangent vector computation (5) which involves an identity transformation and matrix multiplication has time complexity of $O(m^{2}r)$. The rank-1 approximation step uses largest singular value from the SVD of the $m\times r$ tangent vector, and has time complexity of $O(mr^{2})$. Finally, the update rule as shown in (7) which has a time complexity of $O(mr^{2})$. The overall complexity of the algorithm is dominated by the matrix multiplication calculations of time complexity $O(mnr)$. However, unlike GaLore, since we avoid computing SVD operation on the $m\times n$ gradient matrix, which has complexity of $O(nm^{2})$, the overall update step in SubTrack-Grad is still more efficient with respect to time complexity.
Table 5: Time Complexity for SubTrack-Grad Subspace Update
@@
Appendix D Memory and Time Comparison
Table 6 presents the wall-times measured to compare the computational efficiency of SubTrack-Grad and GaLore. Additionally, Table 7 shows the peak memory consumption for each architecture during training, using the hyperparameters detailed in 14, on an NVIDIA A100 GPU with 40GB of memory. These values were used to create the bar plots shown in 2(a). For wall-time comparisons, each architecture was trained for 2,000 steps, ensuring exactly 10 subspace updates while excluding evaluation steps. The number of subspace updates was doubled compared to the fine-tuning experiments because larger models and more complex tasks naturally require more frequent updates. This adjustment allowed for more accurate measurements considering this demand.
Table 6: Wall-time comparison for pre-training Llama-based architectures with varying model sizes on the C4 dataset. The experiments consist of 2000 iterations, corresponding to exactly 10 subspace update steps. The last two columns present the average percentage change in runtime relative to Full-Rank and GaLore, respectively.
60M 130M 350M 1B 3B Avg w.r.t FR w.r.t GaLore
r=128 r=256 r=256 r=512 r=512
Full-Rank 524.0 1035.1 1396.4 974.9 1055.9 997.3 - -
BAdam 511.3 779.2 961.6 798.6 1004.1 811.0 -18.68% -
GaLore 547.8 1094.2 1589.0 1729.5 2715.5 1535.2 +53.94% -
SubTrack-Grad 534.9 1061.0 1465.9 1191.1 1385.7 1127.7 +13.08% -26.54%
Table 7: Peak memory consumption comparison when pre-training Llama-based architectures with different sizes on C4 dataset. The last two columns present the average percentage change in memory consumption relative to Full-Rank and GaLore, respectively.
60M 130M 350M 1B 3B Avg w.r.t FR w.r.t GaLore
r=128 r=256 r=256 r=512 r=512
Full-Rank 16.86 25.32 28.67 18.83 34.92 24.92 - -
BAdam 13.34 20.01 21.11 9.70 15.25 15.88 -36.28% -
GaLore 16.89 25.51 27.85 15.24 26.03 22.30 -10.51% -
SubTrack-Grad 16.89 25.52 27.85 15.24 26.16 22.33 -10.39% +0.13%
Table 8 and Table 9 present the wall-time and peak memory consumption, respectively, for fine-tuning RoBERTa-Base on GLUE tasks. These values were used to generate the bar plots shown in 2(c). For wall-time comparisons, the same settings were used, but evaluation steps were excluded, and fine-tuning was limited to number of iterations to ensure exactly 5 subspace update steps during the process.
Table 8: Wall-time comparison when fine-tuning RoBERTa-Base on GLUE tasks for different ranks $r$ up to a number of iterations to exactly include 5 subspace update steps. The last two columns present the average percentage change in runtime relative to Full-Rank and GaLore, respectively.
COLA STS-B MRPC RTE SST-2 MNLI QNLI QQP Avg w.r.t FR w.r.t GaLore
Full-Rank 105.0 124.0 142.2 260.1 95.0 146.3 161.3 119.0 144.1 - -
BAdam 92.7 103.3 111.4 177.9 85.9 116.0 122.1 99.1 113.5 -21.24% -
GaLore (r=4) 189.2 194.7 198.0 310.7 185.0 205.7 215.1 187.3 210.7 +46.22% -
SubTrack-Grad (r=4) 145.2 155.7 168.3 286.2 141.2 172.6 189.9 151.3 176.3 +22.48% -16.23
GaLore (r=8) 189.9 195.2 195.9 312.3 183.9 205.1 215.7 187.6 210.7 +46.22% -
SubTrack-Grad (r=8) 145.2 156.1 168.4 286.2 143.2 172.0 188.9 152.2 176.6 +22.35 -16.33%
Table 9: Peak memory consumption when fine-tuning RoBERTa-Base on GLUE tasks for different ranks $r$. The last two columns present the average percentage change in memory consumption relative to Full-Rank and GaLore, respectively.
COLA STS-B MRPC RTE SST-2 MNLI QNLI QQP Avg w.r.t FR w.r.t GaLore
Full-Rank 4.13 4.24 4.29 8.16 3.88 7.66 9.50 7.80 6.21 - -
BAdam 2.12 2.15 2.30 3.69 1.99 2.13 2.24 2.84 2.43 -60.87% -
GaLore (r=4) 3.64 4.12 3.85 7.27 3.52 6.84 8.75 6.58 5.57 -10.31% -
SubTrack-Grad (r=4) 3.64 4.12 3.86 7.27 3.37 6.73 8.75 6.58 5.54 -10.79% -0.54%
GaLore (r=8) 3.64 4.12 3.86 7.27 3.53 6.84 8.75 6.58 5.57 -10.31% -
SubTrack-Grad (r=8) 3.65 4.12 3.86 7.27 3.38 6.74 8.75 6.58 5.54 -10.79% -0.54%
Table 10 and Table 11 present the wall-time and peak memory consumption for each method when fine-tuning RoBERTa-Large on SuperGLUE tasks. These values were used to generate the bar plots shown in 2(b). For wall-time comparisons, as with GLUE tasks, evaluation steps were omitted to ensure accurate measurements of fine-tuning time, and the number of iterations was adjusted to include exactly 5 subspace updates during fine-tuning.
Table 10: Wall-time comparison when fine-tuning RoBERTa-Large on SuperGLUEup to a number of iterations to exactly include 5 subspace update steps. The last two columns present the average percentage change in runtime relative to Full-Rank and GaLore, respectively.
BoolQ CB COPA WIC WSC AXg Avg w.r.t FR w.r.t GaLore
Full-Rank 1406.3 171.5 35.1 250.4 216.4 52.1 355.3 - -
BAdam 794.9 93.5 45.9 163.9 137.2 38.1 212.25 -40.26% -
GaLore (r=8) 1536.8 272.2 235.3 483.0 393.2 219.9 523.4 +47.31% -
SubTrack-Grad (r=8) 1468.0 198.1 82.2 326.1 266.7 86.6 404.62 +13.88% -22.69%
Table 11: Peak memory consumption comparison when fine-tuning RoBERTa-Large on SuperGLUE. The last two columns present the average percentage change in memory consumption relative to Full-Rank and GaLore, respectively.
BoolQ CB COPA WIC WSC AXg Avg w.r.t FR w.r.t GaLore
Full-Rank 20.16 15.15 7.86 7.96 9.11 7.81 11.37 - -
BAdam 10.80 5.78 3.36 3.26 4.57 3.26 5.17 -54.52% -
GaLore (r=8) 19.08 12.53 5.32 5.94 7.26 5.34 9.24 -18.73% -
SubTrack-Grad (r=8) 19.09 12.40 5.12 5.95 7.06 5.35 9.16 -19.44% -0.87%
Appendix E Fine-Tuning RoBERTa
RoBERTa-Base was fine-tuned on GLUE tasks using the hyperparameters detailed in Table 12, matching those reported in the GaLore (Zhao et al., 2024) for rank-$4$ and rank-$8$ subspaces, with a subspace update interval set at 500 iterations.

### v2

@@ 6 Discussion and Conclusion
We propose SubTrack++ , a time- and memory-efficient approach that projects gradients into a low-rank subspace and uses Grassmannian subspace tracking to preserve the computed subspace while incorporating gradient components from the orthogonal complement. By integrating projection-aware optimizers that reflect subspace changes in Adam’s internal statistics and utilizing the gradient information lost during low-rank projection, SubTrack++ achieves state-of-the-art convergence and accuracy across all baselines. While Grassmannian subspace tracking integrates seamlessly with various optimizers as a plug-and-play module, extending projection-aware optimization beyond the Adam family requires further design and investigation. Also benchmarking on models with more than 7B parameters was not feasible regarding our limited time and resources.
Appendix A Convergence of SubTrack++
@@
proof. To demonstrate that SubTrack++ converges to the global minimum during training, we begin by deriving the recursive form of the gradients.
@@
While SubTrack++ utilizes right/left projections to reduce memory consumption, the proof is presented using both projection matrices to ensure generality. Here, we demonstrate how the proof proceeds under the assumption $m\leq n$ (without loss of generality), which allows the use of the left projection matrix.
@@ Appendix B Grassmann Exponential
$\Delta=Q\begin{pmatrix}0&B^{\top}\\ B&0\end{pmatrix}Q^{\top}=\left[Q\begin{pmatrix}0&-B^{\top}\\ B&0\end{pmatrix}Q^{\top},P\right]$ (34)
@@
$\Delta^{\text{hor}}_{Q}=[\Delta,P]Q=Q\begin{pmatrix}0&-B^{\top}\\ B&0\end{pmatrix}$ (35)
@@
The orthogonal group $O(n)$ is defined as the set of all $n\times n$ matrices $Q$ over $\mathbb{R}$ such that $Q^{\top}Q=QQ^{\top}=I_{n}$, where $Q^{\top}$ is the transpose of $Q$ and $I_{n}$ is the $n\times n$ identity matrix:
$O(n)=\{Q\in\mathbb{R}^{n\times n}\mid Q^{\top}Q=I_{n}=QQ^{\top}\}.$
Then the tangent space of the orthogonal group $O(n)$ at a point $Q$, denoted $T_{Q}O(n)$, is defined as the set of matrices of the form $Q\Omega$, where $\Omega\in\mathbb{R}^{n\times n}$ is a skew-symmetric matrix, i.e., $\Omega^{\top}=-\Omega$:
$T_{Q}O(n)=\{Q\Omega\mid\Omega\in\mathbb{R}^{n\times n},\Omega^{\top}=-\Omega\}.$
@@
If $P\in Gr(n,p)$ and $\Delta\in T_{P}Gr(n,p)$ with $\Delta^{\text{hor}}_{Q}=Q\begin{pmatrix}0&-B^{\top}\\ B&0\end{pmatrix}$, the geodesic in the Grassmannian is therefore
$\text{Exp}^{Gr}_{P}(t\Delta)=\pi^{OG}\left(Q\exp_{m}\left(t\begin{pmatrix}0&-B^{\top}\\ B&0\end{pmatrix}\right)\right).$ (37)
@@
$B=U_{\perp}^{\top}\Delta^{\text{hor}}_{U}=U_{\perp}^{\top}\hat{Q}\Sigma V^{\top}$
@@
with $W:=U_{\perp}^{\top}\hat{Q}\in St(n-p,r),\Sigma\in\mathbb{R}^{r\times r},V\in St(p,r)$. Let $W_{\perp},V_{\perp}$ be suitable orthogonal completions. Then,
$\exp_{m}\begin{pmatrix}0&-B^{\top}\\ B&0\end{pmatrix}=\begin{pmatrix}V&V_{\perp}&0&0\\ 0&0&W&W_{\perp}\end{pmatrix}\begin{pmatrix}\cos(\Sigma)&0&-\sin(\Sigma)&0\\ 0&I_{p-r}&0&0\\ \sin(\Sigma)&0&\cos(\Sigma)&0\\ 0&0&0&I_{n-p-r}\end{pmatrix}\begin{pmatrix}V^{\top}&0\\ V_{\perp}^{\top}&0\\ 0&W^{\top}\\ 0&W_{\perp}^{\top}\end{pmatrix},$
@@
which leads to the desired result when inserted into (37). For more mathematical details, you can refer to Edelman et al. [1998], Bendokat et al. [2024], or other useful resources on Grassmann geometry.
Appendix C Projection-Aware Optimizer
When projecting into a lower-dimensional space and tracking coordinate changes, we typically use orthonormal projection matrices to represent the subspaces and their multiplications to effect a change of basis. While this works well for purely linear operations, Adam’s updates also incorporate non-linear elements.
Suppose the subspace changes at step $t$, transitioning from the subspace spanned by the orthonormal matrix $S_{t-1}$ to that spanned by $S_{t}$. Since both matrices are orthonormal—preserved by the Grassmannian update rule in (5)—the matrix $S_{t}^{\top}S_{t-1}$ represents the change of basis between the two subspaces. In other words, if $\mathbb{E}_{t-1}=(e_{t-1}^{1},\dots,e_{t-1}^{r})$ and $\mathbb{E}_{t}=(e_{t}^{1},\dots,e_{t}^{r})$ are orthonormal bases for the subspaces at steps $t-1$ and $t$, respectively, then the $i$th column of matrix $X$ transforms under the change of basis as $X^{i}_{t}=\sum_{j=1}^{r}\langle e^{i}_{t},e^{j}_{t-1}\rangle X^{j}_{t-1}$, where $X^{j}_{t-1}$ is the $j$th column of $X$ based on the basis of the time step $t-1$.
$\mathbf{E}_{t,\beta}\left[.\right]$ denotes the exponential time-weighted expectation at time $t$ with decay rate $\beta$. Following the reinterpretation by Robert et al. [2025], Adam’s first and second moment estimates can be expressed as $\widetilde{M}_{t}=\mathbf{E}_{t,\beta_{1}}\left[\widetilde{G}_{t}\right]$ and $\widetilde{\mathcal{V}}_{t}=\mathbf{E}_{t,\beta_{2}}\left[(\widetilde{G}_{t})^{2}\right]$, where $\widetilde{G}_{t}$ denotes the low-rank representation of the gradient at time step $t$. As shown in (38), the first moment estimate can be transformed under a change of basis using the change-of-basis matrix, $S_{t}^{\top}S_{t-1}$. Notably, $\langle\widetilde{G}_{t},e^{i}_{t}\rangle$ gives the $i$th column of $\widetilde{G}_{t}$ when the subspace has the basis $\mathbb{E}_{t}$. We use the superscripts to indicate a column of a matrix.
$\small\mathbf{E}_{t,\beta_{1}}\left[\langle\widetilde{G}_{t},e^{i}_{t}\rangle\right]=\sum_{j=1}^{r}\langle e^{i}_{t},e^{j}_{t-1}\rangle\mathbf{E}_{t,\beta_{1}}\left[\langle\widetilde{G}_{t},e^{j}_{t-1}\rangle\right]=\sum_{j=1}^{r}\langle e^{i}_{t},e^{j}_{t-1}\rangle\widetilde{M}_{t}^{j}=\left(S_{t}^{\top}S_{t-1}\widetilde{M}_{t}\right)^{i}$ (38)
Following the same approach, we can change the basis for the second moment estimate as described in (39).
$\small\mathbf{E}_{t,\beta_{2}}\left[\left(\langle\widetilde{G}_{t},e^{i}_{t}\rangle\right)^{2}\right]$ $=\sum_{j=1}^{r}\langle e^{i}_{t},e^{j}_{t-1}\rangle^{2}\,\mathbf{E}_{t,\beta_{2}}\left[\left(\langle\widetilde{G}_{t},e^{j}_{t-1}\rangle\right)^{2}\right]$ (39)
$+\sum_{k\neq l}^{r}\langle e^{i}_{t},e^{k}_{t-1}\rangle\langle e^{i}_{t},e^{l}_{t-1}\rangle\,\mathbf{E}_{t,\beta_{2}}\left[\langle\widetilde{G}_{t},e^{k}_{t-1}\rangle\langle\widetilde{G}_{t},e^{l}_{t-1}\rangle\right]$
$=\sum_{j=1}^{r}\langle e^{i}_{t},e^{j}_{t-1}\rangle^{2}\widetilde{\mathcal{V}}_{t}^{j}$
$+\sum_{k\neq l}^{r}\langle e^{i}_{t},e^{k}_{t-1}\rangle\langle e^{i}_{t},e^{l}_{t-1}\rangle\widetilde{M}_{t}^{k}\widetilde{M}_{t}^{l}.$
In transitioning from the first equality to the second, we assume independence among the gradient coordinates. This enables us to approximate the covariance using a product of first-order moment estimates. This assumption is often reasonable in practice because we compute the SVD of the gradient and maintain an orthonormal subspace projection matrix, updating it along the Grassmannian geodesic to track the optimal subspace. Since SVD tends to diagonalize the covariance, the off-diagonal entries are typically negligible. Additionally, we clip any negative values to zero to ensure valid (non-negative) variance estimates. Moreover, to rewrite the second term in the final equality of (39), we employ the following equation:
$\sum_{k}\sum_{l}\langle e^{i}_{t},e^{k}_{t-1}\rangle\langle e^{i}_{t},e^{l}_{t-1}\rangle\widetilde{M}_{t}^{k}\widetilde{M}_{t}^{l}$ (40)
$=\sum_{k}\langle e^{i}_{t},e^{k}_{t-1}\rangle^{2}\left(\widetilde{M}_{t}^{k}\right)^{2}\;+\;\sum_{k\neq l}\langle e^{i}_{t},e^{k}_{t-1}\rangle\langle e^{i}_{t},e^{l}_{t-1}\rangle\widetilde{M}_{t}^{k}\widetilde{M}_{t}^{l}$
Given these, we can rewrite (39) as follows:
$\mathbf{E}_{t,\beta_{2}}\left[\left\langle\widetilde{G}_{t},e_{t}^{i}\right\rangle^{2}\right]$ $=\sum_{j}\langle e^{i}_{t},e^{j}_{t-1}\rangle^{2}\widetilde{\mathcal{V}}_{t}^{j}+$ (41)
$\left[\sum_{k}\sum_{l}\langle e^{i}_{t},e^{k}_{t-1}\rangle\langle e^{i}_{t},e^{l}_{t-1}\rangle\widetilde{M}_{t}^{k}\widetilde{M}_{t}^{l}-\sum_{k}\langle e^{i}_{t},e^{k}_{t-1}\rangle^{2}\left(\widetilde{M}_{t}^{k}\right)^{2}\right]$
$=\sum_{j}\langle e^{i}_{t},e^{j}_{t-1}\rangle^{2}\left[\widetilde{\mathcal{V}}_{t}^{j}-\left(\widetilde{M}_{t}^{j}\right)^{2}\right]+\left(\langle e^{i}_{t},e^{j}_{t-1}\rangle\widetilde{M}_{t}{}\right)^{2}$
$=\left(\left(S_{t}^{\top}S_{t-1}\right)^{2}\left[\widetilde{\mathcal{V}}_{t}-\widetilde{M}_{t}^{2}\right]\right)^{i}+\left(\left(S_{t}^{\top}S_{t-1}\widetilde{M}_{t}\right)^{2}\right)^{i}$
By applying (38) and (41), we can directly derive the update rules for the projection-aware optimizer, as expressed in (8) and (9).
Appendix D Time Complexity Analysis
@@
Table 3 presents the time complexity breakdown for the subspace update step in the SubTrack++ algorithm assuming a $m\times n$ gradient matrix and rank $r$ projection matrix, where $r\ll m\leq n$. As outlined in Algorithm 1, the subspace update step begins by solving the least squares problem (2) to estimate the optimal update for $S_{t}$, the $m\times r$ orthonormal matrix. This operation has a time complexity of $O(mr^{2})$. Computing the residual and the partial derivative with respect to $S_{t}$ requires $O(mrn)$ and $O(mnr)$ time respectively. This is because the solution to the least squares problem, $A$, has shape $r\times n$ which is multiplied by $S_{t}$ in the residual $R=G_{t}-S_{t}A$, resulting in time complexity $O(mrn)$. The following operation for the partial derivative is $-2RA^{T}$, where the matrix multiplication has $O(mnr)$ complexity. The tangent vector computation (4) which involves an identity transformation and matrix multiplication has time complexity of $O(m^{2}r)$.
Table 3: Time Complexity for SubTrack++ Subspace Update
@@
The rank-1 approximation step uses largest singular value from the SVD of the $m\times r$ tangent vector, and has time complexity of $O(mr^{2})$. Finally, the update rule as shown in (13) which has a time complexity of $O(mr^{2})$. The overall complexity of the algorithm is dominated by the matrix multiplication calculations of time complexity $O(mnr)$. However, unlike GaLore, since we avoid computing SVD operation on the $m\times n$ gradient matrix, which has complexity of $O(nm^{2})$, the overall update step in SubTrack++ is still more efficient with respect to time complexity.
Appendix E Fine-Tuning Experiments
As described, we examined SubTrack++ on fine-tuning RoBERTa-base and RoBERTa-large models to evaluate them on GLUE and SuperGLUE benchmarks. The results for GLUE task is summarized in Table 4, and SuperGLUE in 5.
Table 4: Evaluating the performance of SubTrack++ and other baselines when fine-tuning RoBERTa-Base on GLUE tasks for $r=8$. The performance is measured via Accuracy for SST-2 and RTE tasks, F1 for MRPC, Pearson Correlation for STS-B, and Matthews Correlation for COLA. The best results are marked in bold, with the second-best performance underlined.
COLA STS-B MRPC RTE SST-2
Full-Rank 62.57 91.03 91.32 77.98 94.27
BAdam [Luo et al., 2024] 54.44 89.01 91.35 68.59 94.15
GaLore [Zhao et al., 2024] 58.54 90.61 91.30 74.37 94.50
LDAdam [Robert et al., 2025] 58.81 90.90 92.22 76.53 94.27
SubTrack++ (Ours) 58.29 90.90 91.84 76.53 90.83
Table 5: Evaluating the performance of SubTrack++ and other baselines when fine-tuning RoBERTa-Large on SuperGLUE tasks with $r=8$. The performance is measured via Accuracy for COPA, WIC, WSC, BoolQ, and AXg tasks, and F1 for CB. The best results are marked in bold, with the second-best performance underlined.
BoolQ CB COPA WIC WSC AXg
Full-Rank 85.96 90.33 76.00 71.79 63.46 96.30
GaLore [Zhao et al., 2024] 85.44 88.85 80.00 71.47 63.46 100.00
BAdam [Luo et al., 2024] 82.51 53.28 59.00 70.38 60.58 51.85
LDAdam Robert et al. [2025] 85.75 56.16 80.00 71.00 64.42 70.37
SubTrack++ (Ours) 84.68 89.66 88.00 70.06 63.46 100.00